OpenStudy (anonymous):

Use an Addition or Subtraction Formula to find the exact value of the expression, as demonstrated in Example 1. cos 165°

OpenStudy (therealmeeeee):

sin(A+B) = sin(A)cos(B)+cos(A)sin(B) tan(A+B) = [tan(A)+tan(B)]/[1-tan(A)tan(B)] sin(45degree) = 1/[2^(1/2)] = [2^(1/2)]/2 cos(45degree) = 1/[2^(1/2)] = [2^(1/2)]/2 tan(45degree) = 1 sin(120degree) = [3^(1/2)]/2 cos(120degree) = -1/2 tan(120degree) = -[3^(1/2)] sin(165) = sin(45+120) = sin(45)cos(120)+cos(45)sin(120) sin(165) = [2^(1/2)]/2*(-1/2) + [2^(1/2)]/2*[3^(1/2)]/2 = -[2^(1/2)]/4+[3^(1/2)]/4 sin(165) = [sqrt(3)-sqrt(2)]/4 tan(165) = tan(45+120) = [tan(45)+tan(120)/[1-tan(45)tan(120)] tan(165) = [1+(-3^(1/2)]/[1-1[-3^(1/2)] tan(165) = (1-sqrt(3))/(1+sqrt(3))

OpenStudy (therealmeeeee):

do this help

OpenStudy (anonymous):

No

OpenStudy (anonymous):

): sorry

OpenStudy (anonymous):

Help please